How much work does the force you apply do on the car?

You apply a constant force F=(-68.0N)i+(36.0N)j to a 420 kg car as the car travels 47.0 m in a direction that is 240.0 counterclockwise from the +-axis.
Answer #1

we have the displacement vector = (41*cos240)i + (41*sin240)j = -20.5i + -35.5j
work done =

= +

= -68*(-20.5) + 36*(-35.5) = 116 joules

so,the work done by the force = 116 joules